# Quantifying Bursty Star Formation in Dwarf Galaxies

### v1

@@ 1 Introduction
Standard galaxy formation in the $\Lambda$CDM paradigm suggests that dwarf galaxies (stellar masses $\sim 10^{3}-10^{9}~{}{\rm M}_{\odot}$) inherently have bursty or episodic star formation histories. Due to their low masses, stellar feedback can temporarily quench star formation, causing the instantaneous star formation rate to fluctuate by orders of magnitude over 10-100 Myr timescales (e.g., Governato et al., 2010; Hopkins et al., 2014; Chan et al., 2015; Dutton et al., 2016; Tollet et al., 2016; Bose et al., 2019; Emami et al., 2019). In the faintest dwarf galaxies, supernova feedback can completely quench star formation for hundreds of Myr (e.g., Bland-Hawthorn et al., 2015; Wheeler et al., 2019) or even permanently (Gallart et al., 2021; Chen et al., 2022). Reionization also appears to force many faint dwarf galaxies to pause star formation for a few Gyr before resuming (Savino et al., 2023; McQuinn et al., 2023, 2024).
@@
Such bursty star formation is crucial for reconciling observed dwarf galaxy central density profiles with $\Lambda$CDM predictions. It is this burstiness that causes initially cuspy Navarro et al. (1996) halos to transform into the observed cored density profiles (e.g., Pontzen & Governato, 2012; Chan et al., 2015; Tollet et al., 2016). Furthermore, JWST observations suggest that dwarf galaxies may over-ionize the IGM, with bursty star formation proposed as a possible solution (Sun et al., 2023; Clarke et al., 2024; Muñoz et al., 2024; Simmonds et al., 2024). Burstiness has also been linked to the miniquenching of high-redshift low-mass galaxies (Antwi-Danso, 2024; Dome et al., 2024; Looser et al., 2024).
@@
In principle, bursty star formation could imprint upon a dwarf galaxy’s chemical abundance distribution, such as the pattern of [$\alpha$/Fe] vs [Fe/H] of individually resolved stars and the stellar metallicity distribution function. Fe is produced by both core-collapse supernovae (CCSNe) and Type Ia supernovae (SNeIa), while $\alpha$-elements are produced primarily by CCSNe. A galaxy experiencing a significant quiescent period will cease $\alpha$-element production, while Fe production by SNeIa continues. Consequently, when star formation resumes, there could be discrete jumps in the [Fe/H] distribution of newly formed stars. These jumps would be visible in sufficiently populated chemical abundance distributions. Such features might be detectable in current dwarf galaxy data samples obtained over the past decade (e.g., Cohen & Huang, 2010; Kirby et al., 2011a; Starkenburg et al., 2013; Hendricks et al., 2014; Hill et al., 2019; Kirby et al., 2019, 2020; Hasselquist et al., 2021), as well as in future surveys (e.g., Skúladóttir et al., 2023).
@@
The majority of chemical evolution models employ smooth prescriptions for gas and star formation histories (SFH) (e.g., Cescutti, 2008; Kirby et al., 2011a; Pilkington & Gibson, 2012; Vincenzo et al., 2014; Recchi & Kroupa, 2015; Ishimaru et al., 2015; Yuan et al., 2016; Côté et al., 2017; Weinberg et al., 2017; Matteucci, 2021; de los Reyes et al., 2022; Johnson et al., 2023; Sandford et al., 2024), with some notable exceptions (e.g., Hirashita et al., 2001; Johnson & Weinberg, 2020). Some semi-analytic models, based on dark matter merger trees, allow for episodic gas accretion, but do not study the impact on individual metallicity distributions (e.g., Lu et al., 2011; Benson et al., 2012; Lee et al., 2014; Benson et al., 2016; Qin et al., 2018; Manwadkar & Kravtsov, 2022). Cosmological simulations tracking chemistry naturally include bursty star formation (e.g., Jeon et al., 2017; Wheeler et al., 2019; Applebaum et al., 2021), but they are still relatively limited in mass and time resolution for studying the early chemical evolution of dwarf galaxies and cannot directly model individual known galaxies.
@@ 2 Simple Chemical Evolution Model
$D_{\text{Ia}}(t)=\begin{dcases}0&t<100\text{ Myr}\\ 0.429\left(\frac{t}{\text{Myr}}\right)^{-1.1}&t\geq 100\text{ Myr}\end{dcases}$ (2)
@@ 4 Measuring the Discontinuous Chemical Track
where $\boldsymbol{\theta}=\{\pi_{k},\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}\}_{k=1}^{K}$ represents the set of all parameters in the GMM, and $N$ is the number of data points. We fit GMMs with 100 random initializations to ensure good convergence.
@@
Based on the AIC score, the model will determine the optimal number of Gaussian components that best describe the data. If the best-fitting model has multiple components, it suggests the presence of multimodality in the chemical space, which would be indicative of bursty star formation.
To demonstrate the effectiveness of GMMs in detecting discontinuous chemical tracks, we start with a mock sample as depicted in Figure 1. We consider two scenarios: the fiducial model with regular starbursts spaced between 150-450 Myr apart (the blue SFH in the top right panel of Figure 1) with each starburst lasting 150 Myr; and the continuous smooth star formation model that integrates to the same total stellar mass of $10^{7}{\rm M}_{\odot}$ (the dashed orange line in Figure 1). We draw samples according to the respective star formation rates and assume an uncertainty of 0.05 dex in both [Fe/H] and [Mg/Fe], similar to the quoted precision and intrinsic dispersion of APOGEE abundances in Sculptor (Mead et al., 2024)
@@
This situation improves if we can further extend the sample size to 1000 stars, as shown in the middle panel. In this case, the sampling noise on the AIC scores decrease substantially and improve the ability to detect multimodality. Obtaining abundances of over 1000 stars in massive dSphs like Sagittarius, Fornax, and Sculptor will soon be possible with spectroscopic surveys on 4-8m telescopes like 4DWARFS (Skúladóttir et al., 2023) and Subaru PFS (Takada et al., 2014). The advent of ELTs will enable such measurements in all but the faintest dSph galaxies (Ji et al., 2019).
@@ 5 Dectecting Bursty Star Formation in Sculptor dSph
To more robustly quantify the multimodality, we calculate the AIC scoree estimate the uncertainty of the AIC score by bootstrapping the data. The AIC score as a function of the number of components is shown in the blue band in the right panel of Figure 2. Just like in the case of the simulated data in the left panel, the real data also exhibits a similar trend, with the AIC score decreasing as the number of components increases, signaling a multimodal nature in the [Mg/Fe]-[Fe/H] plane. For comparison, we show the AIC score for a smoothed data example as a green band in the right panel. We generate the smoothed data by fitting the real data with a two-component GMM [^1] We choose two components because there is an apparent kink in the chemical track around [Fe/H] = -1.3 (see top panel of Figure 3), but the results are practically the same if we were to fit a single Gaussian distribution., draw from the distribution, and ensure that the smoothed distribution follows the contours of the data. In this case, the smoothed data shows no trend in the AIC scores, demonstrating that the real data has a statistical tendency to favor the case of multimodality.
@@ 6.2 Intrinsic spread in [$\alpha$/Fe] from the starbursts
Initially, we had anticipated that the [Mg/Fe] spread would be an independent useful indicator of bursty star formation. This was motivated by past studies (e.g., Recchi & Kroupa, 2015; Muley et al., 2021; Patel et al., 2022; Mead et al., 2024) and by the sharp sawtooth [Mg/Fe] vs [Fe/H] pattern in Figure 1. However, detecting and using the [Mg/Fe] spread turns out to be substantially more difficult than the [Fe/H] gaps in the model presented here. Observationally, this is because it is much harder to detect an intrinsic scatter than a difference in means: the uncertainty on intrinsic scatter scales as $N_{*}^{1/4}$ while means scales as $N_{*}^{1/2}$ (see Ting & Weinberg, 2022). Theoretically, it is because star formation bursts have to be somewhat fine-tuned to capture the maximum [Mg/Fe] spread: a burst must be longer than 10 Myr to capture any Mg being produced, and shorter than 100 Myr so Type Ia supernovae do not start contributing to Fe. Even when this happens, the signal is quite small.
@@ 6.4 Starburst Signals from Other Dwarf Galaxies
Table 1: References. 1: McQuinn et al. 2024, 2: Norris et al. 2017a, 3: Norris et al. 2017b, 4: Simon et al. 2023, 5: Ji et al. 2023 and Luna et al. in prep.
@@ 6.5 Caveats and Future Directions
In this study, we use a constant effective mixing gas mass $M_{\rm g,eff}$. While this simplifies the description of the metallicity gap, it does so at the cost of making the gas mass less interpretable: $M_{\rm g,eff}$ is a degenerate combination of gas and metal inflows and outflows. In contrast, standard chemical evolution models tie the gas mass (usually the cold gas within the galaxy) to the star formation and outflow rates through a star formation efficiency and mass loading factor or prescription (e.g., Cescutti, 2008; Salvadori et al., 2008; Ishimaru et al., 2015; Weinberg et al., 2017; Johnson et al., 2023). While such parameterizations are physically well-motivated and provide good results for overall dwarf galaxy scaling relations, current implementations prohibit the type of bursty star formation seen in hydrodynamic simulations. We will explore the interpretation and degeneracies of $M_{\rm g,eff}$ in more detail in future work.

### v2

@@ 1 Introduction
Standard galaxy formation in the $\Lambda$CDM paradigm suggests that dwarf galaxies (stellar masses $\sim 10^{3}-10^{9}~{}{\rm M}_{\odot}$) inherently have bursty or episodic star formation histories. Due to their low masses, stellar feedback can temporarily quench star formation, causing the instantaneous star formation rate to fluctuate by orders of magnitude over 10-100 Myr timescales (e.g., Governato et al., 2010; Hopkins et al., 2014; Chan et al., 2015; Dutton et al., 2016; Tollet et al., 2016; Bose et al., 2019; Emami et al., 2019). In the faintest dwarf galaxies, supernova feedback can completely quench star formation for hundreds of Myr (e.g., Bland-Hawthorn et al., 2015; Wheeler et al., 2019; Hirai et al., 2024) or even permanently (Gallart et al., 2021; Chen et al., 2022). Reionization also appears to force many faint dwarf galaxies to pause star formation for a few Gyr before resuming (Savino et al., 2023; McQuinn et al., 2023, 2024).
@@
Such bursty star formation is crucial for reconciling observed dwarf galaxy central density profiles with $\Lambda$CDM predictions. It is this burstiness that causes initially cuspy Navarro et al. (1996) halos to transform into the observed cored density profiles (e.g., Pontzen and Governato, 2012; Chan et al., 2015; Tollet et al., 2016). Furthermore, JWST observations suggest that dwarf galaxies may over-ionize the IGM, with bursty star formation proposed as a possible solution (Sun et al., 2023; Clarke et al., 2024; Muñoz et al., 2024; Simmonds et al., 2024). Burstiness has also been linked to the miniquenching of high-redshift low-mass galaxies (Antwi-Danso, 2024; Dome et al., 2024; Looser et al., 2024).
@@
In principle, bursty star formation could imprint upon a dwarf galaxy’s chemical abundance distribution, such as the pattern of [$\alpha$/Fe] vs [Fe/H] of individually resolved stars and the stellar metallicity distribution function. Fe is produced by both core-collapse supernovae (CCSNe) and Type Ia supernovae (SNeIa), while $\alpha$-elements are produced primarily by CCSNe. A galaxy experiencing a significant quiescent period will cease $\alpha$-element production, while Fe production by SNeIa continues. Consequently, when star formation resumes, there could be discrete jumps in the [Fe/H] distribution of newly formed stars. These jumps would be visible in sufficiently populated chemical abundance distributions. Such features might be detectable in current dwarf galaxy data samples obtained over the past decade (e.g., Cohen and Huang, 2010; Kirby et al., 2011a; Starkenburg et al., 2013; Hendricks et al., 2014; Hill et al., 2019; Kirby et al., 2019, 2020; Hasselquist et al., 2021), as well as in future surveys (e.g., Skúladóttir et al., 2023).
@@
The majority of chemical evolution models employ smooth prescriptions for gas and star formation histories (SFH) (e.g., Cescutti, 2008; Kirby et al., 2011a; Pilkington and Gibson, 2012; Vincenzo et al., 2014; Recchi and Kroupa, 2015; Ishimaru et al., 2015; Yuan et al., 2016; Côté et al., 2017; Weinberg et al., 2017; Matteucci, 2021; de los Reyes et al., 2022; Johnson et al., 2023; Sandford et al., 2024), with some notable exceptions (e.g., Hirashita et al., 2001; Johnson and Weinberg, 2020). Some semi-analytic models, based on dark matter merger trees, allow for episodic gas accretion, but do not study the impact on individual metallicity distributions (e.g., Lu et al., 2011; Benson et al., 2012; Lee et al., 2014; Benson et al., 2016; Qin et al., 2018; Manwadkar and Kravtsov, 2022). Cosmological simulations tracking chemistry naturally include bursty star formation (e.g., Jeon et al., 2017; Wheeler et al., 2019; Applebaum et al., 2021), but they are still relatively limited in mass and time resolution for studying the early chemical evolution of dwarf galaxies and cannot directly model individual known galaxies.
@@ 2 Simple Chemical Evolution Model
$D_{\text{Ia}}(t)=\begin{cases}0&\text{for }t<100\text{ Myr}\\ 0.429\left(\frac{t}{\text{Myr}}\right)^{-1.1}&\text{for }t\geq 100\text{ Myr}\end{cases}$ (2)
@@
Our choice of a semi-analytical framework, while simplified compared to hydrodynamical simulations, provides two key advantages: analytical tractability that allows us to directly connect observable features to physical parameters, and the ability to isolate the specific impact of bursty star formation on chemical evolution. This approach complements existing hydrodynamical simulations that naturally include bursty star formation but cannot yet necessarily achieve the mass and time resolution needed for studying the early chemical evolution of individual known dwarf galaxies (Jeon et al., 2017; Wheeler et al., 2019; Hirai et al., 2024).
@@ 4 Measuring the Discontinuous Chemical Track
GMMs are particularly well-suited for our analysis for several reasons. The strong inductive bias - which assumes the data can be modeled as a mixture of Gaussian components - allows us to perform robust modeling with our small dataset ($\sim$100 stars). While such assumption is certainly a limitation, in the absence of measurement uncertainties, chemical evolution tracks follow well-defined one-dimensional paths, with observed spread primarily reflecting measurement uncertainties. And chemical abundance uncertainties are approximately Gaussian, as they derive from combining multiple spectral features. We note that while more flexible density estimation methods exist (e.g., Normalizing Flows Ting and Weinberg, 2022), GMMs explicitly model discrete subcomponents - exactly what we expect from episodic star formation. As we will see, to ensure robust results despite these limitations, we carefully validate our GMM fits through bootstrap resampling and comparison with smoothed models (see Figure 2) with simulated data, finding the method is adequate to find statistical evidence for multimodality in our dataset.
@@
where $\boldsymbol{\theta}=\{\pi_{k},\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}\}_{k=1}^{K}$ represents the set of all parameters in the GMM, and $N$ is the number of data points.
Finding the optimal parameters for a GMM is a complex non-convex optimization problem that cannot be solved analytically. We therefore fit GMMs using the Expectation-Maximization (EM) algorithm implemented in scikit-learn, which iteratively estimates the model parameters by alternating between two steps: (1) computing the probability of each data point belonging to each component (E-step), and (2) updating the model parameters to maximize the likelihood given these probabilities (M-step). The EM algorithm is guaranteed to converge to a local optimum, but may not find the global optimum. To address this limitation, we perform 1000 independent fits with different random initializations, validating that each component has at least 4 points and covariance correlations below 0.9, and select the model with the highest likelihood score. This approach helps avoid poor local optima and ensures robust convergence to a well-behaved solution that accurately captures the underlying structure in the data.
@@
Based on the AIC score, we determine the optimal number of Gaussian components that best describe the data. The AIC balances model fit against model complexity according to the equation above. A lower AIC indicates a more favorable trade-off between complexity and fit. Specifically, $-2\ln(\hat{L})$ rewards models that better fit the data, while $2m$ penalizes complex models. Under the assumption that the true model lies within the family of candidate models, AIC is asymptotically optimal in selecting the model closest (in a Kullback–Leibler sense) to the true process (Akaike, 1974). When the preferred model includes multiple Gaussian components, this naturally suggests multimodality in the chemical abundance space, pointing to multiple subpopulations. In an astrophysical context, such a pattern is consistent with bursty star formation, wherein different episodes of star formation give rise to distinct chemical loci.
To demonstrate the effectiveness of GMMs in detecting discontinuous chemical tracks, we start with a mock sample as depicted in Figure 1. We consider two scenarios: the fiducial model with regular starbursts spaced between 150-450 Myr apart (the blue SFH in the top right panel of Figure 1) with each starburst lasting 150 Myr; and the continuous smooth star formation model that integrates to the same total stellar mass of $10^{7}{\rm M}_{\odot}$ (the dashed orange line in Figure 1). We draw samples according to the respective star formation rates and assume an uncertainty of 0.05 dex in both [Fe/H] and [Mg/Fe], similar to the quoted precision and intrinsic dispersion of APOGEE abundances in Sculptor (Mead et al., 2024).
@@
This situation improves if we can further extend the sample size to 1000 stars, as shown in the middle panel. In this case, the sampling noise on the AIC scores decrease substantially and improve the ability to detect multimodality. Obtaining abundances of over 1000 stars in massive dSphs like Sagittarius, Fornax, and Sculptor will soon be possible with spectroscopic surveys on 4-8m telescopes like 4DWARFS (Skúladóttir et al., 2023) and Subaru PFS (Takada et al., 2014). The advent of ELTs will enable such measurements in all but the faintest dSph galaxies (Ji et al., 2019). Although not shown, our tests indicate that measurement precision is less critical than sample size for detecting multimodality in these systems. The metallicity gaps, particularly at low [Fe/H], are typically larger than 0.05 dex, exceeding typical spectroscopic uncertainties. However, at higher metallicities ([Fe/H] $>-1.5$) and for shorter starburst gaps ($\sim$100 Myr), higher precision measurements could help better resolve the multimodal structure (see Section 6)
@@ 5 Dectecting Bursty Star Formation in Sculptor dSph
To more robustly quantify the multimodality, we calculate the AIC score estimate the uncertainty of the AIC score by bootstrapping the data. The AIC score as a function of the number of components is shown in the blue band in the right panel of Figure 2. Just like in the case of the simulated data in the left panel, the real data also exhibits a similar trend, with the AIC score decreasing as the number of components increases, signaling a multimodal nature in the [Mg/Fe]-[Fe/H] plane. For comparison, we show the AIC score for a smoothed data example as a green band in the right panel. We generate the smoothed data by fitting the real data with a two-component GMM [^1] We choose two components because there is an apparent kink in the chemical track around [Fe/H] = -1.3 (see top panel of Figure 3), but the results are practically the same if we were to fit a single Gaussian distribution., draw from the distribution, and ensure that the smoothed distribution follows the contours of the data. In this case, the smoothed data shows no trend in the AIC scores, demonstrating that the real data has a statistical tendency to favor the case of multimodality. The close agreement between the AIC behavior of our simulated bursty star formation models and the actual Sculptor data provides statistical evidence linking the observed multimodality to discrete star formation events, particularly given the clear contrast with the flat AIC scores seen in our continuous star formation simulations.
@@ 6.2 Intrinsic spread in [$\alpha$/Fe] from the starbursts
Initially, we had anticipated that the [Mg/Fe] spread would be an independent useful indicator of bursty star formation. This was motivated by past studies (e.g., Recchi and Kroupa, 2015; Muley et al., 2021; Patel et al., 2022; Mead et al., 2024) and by the sharp sawtooth [Mg/Fe] vs [Fe/H] pattern in Figure 1. However, detecting and using the [Mg/Fe] spread turns out to be substantially more difficult than the [Fe/H] gaps in the model presented here. Observationally, this is because it is much harder to detect an intrinsic scatter than a difference in means: the uncertainty on intrinsic scatter scales as $N_{*}^{1/4}$ while means scales as $N_{*}^{1/2}$ (see Ting and Weinberg, 2022). Theoretically, it is because star formation bursts have to be somewhat fine-tuned to capture the maximum [Mg/Fe] spread: a burst must be longer than 10 Myr to capture any Mg being produced, and shorter than 100 Myr so Type Ia supernovae do not start contributing to Fe. Even when this happens, the signal is quite small.
@@ 6.4 Starburst Signals from Other Dwarf Galaxies
Table 1: Comparison of star formation quiescent periods ($\delta t_{\rm gap}$) and metallicity gaps in dwarf galaxies. For each galaxy, we list the measured time gap between star formation episodes, the metallicity at which the gap is measured, the observed metallicity gap in the [Fe/H] distribution, and our model’s predicted gap size.
References. 1: McQuinn et al. 2024, 2: Norris et al. 2017a, 3: Norris et al. 2017b, 4: Simon et al. 2023, 5: Ji et al. 2023 and Luna et al. in prep.
@@ 6.5 Caveats and Future Directions
In this study, we use a constant effective mixing gas mass $M_{\rm g,eff}$. While this simplifies the description of the metallicity gap, it does so at the cost of making the gas mass less interpretable: $M_{\rm g,eff}$ is a degenerate combination of gas and metal inflows and outflows. In contrast, standard chemical evolution models tie the gas mass (usually the cold gas within the galaxy) to the star formation and outflow rates through a star formation efficiency and mass loading factor or prescription (e.g., Cescutti, 2008; Salvadori et al., 2008; Ishimaru et al., 2015; Weinberg et al., 2017; Johnson et al., 2023). While such parameterizations are physically well-motivated and provide good results for overall dwarf galaxy scaling relations, current implementations prohibit the type of bursty star formation seen in hydrodynamic simulations.
The effectiveness of our constant $M_{\rm g,eff}$ assumption likely varies with galaxy mass. More massive dwarf galaxies with deeper potential wells may maintain more stable gas reservoirs, making our assumption more applicable. Conversely, ultra-faint dwarfs with stronger feedback effects might require more sophisticated gas mass evolution models. However, our successful reproduction of Sculptor’s chemical patterns suggests that for intermediate-mass dwarf spheroidals, the constant $M_{\rm g,eff}$ approximation captures the essential physics needed to interpret chemical discontinuities as markers of bursty star formation. Future work incorporating time-varying gas masses will help quantify how feedback efficiency impacts the visibility and interpretation of these chemical signatures across different mass scales (e.g., Agertz et al., 2013; Read et al., 2016).